# Exclusively Penalized Q-learning for Offline Reinforcement Learning

### v1

@@ 4 Experiments
4.1 Performance Comparisons
We compare our algorithm with various constraint-based offline RL methods, including CQL baselines [15] on which our algorithm is based on. For other baseline methods, we consider behavior cloning (BC) and 10% BC, where the latter only utilizes only the top 10% of demonstrations with high returns, TD3+BC [21] that simply combines BC with TD3 [5], Onestep RL [22] that performs a single policy iteration based on the dataset, implicit $Q$-learning (IQL) [23] that seeks the optimal value function for the dataset through expectile regression, mildly conservative $Q$-learning (MCQ) [24] that reduces overestimation by using pseudo $Q$ values for out-of-distribution actions, and MISA [25] that considers the policy constraint based on mutual information. To assess baseline algorithm performance, we utilize results directly from the original papers for CQL (paper) [15] and MCQ [24], as well as reported results from other baseline algorithms according to Ma et al. [25]. For CQL, reproducing its performance is challenging, so we also include reproduced CQL performance labeled as CQL (reprod.) from Ma et al. [25]. Any missing experimental results have been filled in by re-implementing each baseline algorithm. For our algorithm, we explored various penalty control thresholds $\tau\in\{c\cdot\rho,~{}c\in[0,10]\}$, where $\rho$ represents the density of $\textrm{Unif}(\mathcal{A})$. For Mujoco tasks, the EPQ penalizing constant is fixed at $\alpha=20.0$, and for Adroit tasks, we consider either $\alpha=5.0$ or $\alpha=20.0$. To ensure robustness, we run our algorithm with four different seeds for each task. Table 1 displays the average normalized returns and corresponding standard deviations for compared algorithms. The performance of EPQ is based on the best hyperparameter setup, with additional results presented in the ablation study in Section 4.3. Further details on the hyperparameter setup are provided in Appendix C.
@@ 4.1 Performance Comparisons
4.2 The Analysis of Estimation Bias
In Section 4.1, EPQ outperforms CQL baselines significantly across various D4RL tasks based on our proposed penalty in Section 3. To analyze the impact of our overestimation reduction method on performance enhancement, we compare the estimation bias for EPQ and CQL baselines with various penalizing constants $\alpha\in\{0,1,5,10,20\}$ on ‘Hopper-random’, ‘Hopper-medium’, and ‘Halfcheetah-medium’ tasks. In Fig. 5(a)(a), we depict the squared value of estimation bias, obtained from the difference between the $Q$-value and the empirical average return for sample trajectories generated by the policy, to show both overestimation bias and underestimation bias. Fig. 5(a)(b) illustrates the corresponding normalized average returns, emphasizing learning progress after $200k$ gradient steps.
In Fig. 5(a)(a), we observe an increase in estimation bias for CQL as the penalizing constant $\alpha$ rises, attributed to unnecessary bias highlighted in Fig. 1. Reducing $\alpha$ to nearly 0 in CQL, however, fails to effectively mitigate overestimation error, leading to a divergence of the $Q$-function in tasks such as ‘Hopper-random’ and ‘Hopper-medium’, as shown in Fig. 1. Conversely, EPQ demonstrates superior reduction of estimation bias in the $Q$-function compared to CQL baselines for all tasks in Fig. 5(a)(a), indicating its capability to mitigate both overestimation and underestimation bias based on the proposed penalty. As a result, Fig. 5(a)(b) shows that EPQ significantly outperforms all CQL variants on ‘Hopper-random’, ‘Hopper-medium’, and ‘Halfcheetah-medium’ tasks.
@@ 4.3 Ablation Study
To understand the impact of EPQ’s components and hyperparameters, we conduct ablation studies to evaluate each component and the penalty control threshold $\tau$ on the ’Hopper-medium’ task.
Component Evaluation: In Section 3, we introduced two variants of the EPQ algorithm: EPQ (w/o PD), which does not incorporate a prioritized dataset as in equation (3), and EPQ (with PD), which leverages a prioritized dataset based on $\hat{\beta}^{Q}$ as in equation (4). In Fig. LABEL:fig:ablation(a), we compare the performance of EPQ (w/o PD), EPQ (with PD), and the CQL baseline to analyze the impact of each component. EPQ (w/o PD) still outperforms CQL, demonstrating that the proposed penalty $\mathcal{P}_{\tau}$ in Section 3.2 enhances performance by efficiently reducing overestimation without introducing unnecessary estimation bias, as discussed in Section 3.2. Additionally, Fig. LABEL:fig:ablation shows that EPQ (with PD) outperforms EPQ (w/o PD), indicating that the proposed prioritized dataset contributes to improved performance, as anticipated in Section 3.3.
Penalty Control Threshold $\tau$: As discussed in Section 3.2, EPQ can dynamically control the penalty amount based on the penalty control threshold $\tau$, as illustrated in Fig. 2, even in the absence of knowledge about the exact number of data samples. Fig. LABEL:fig:ablation(b) demonstrates the performance of EPQ on ‘Hopper-medium’ with various penalty control thresholds $\tau$. From the result, $\tau=0.2\rho$ yields the best performance, and surprisingly, the performance appears to be less sensitive to changes in $\tau$. We initially expect that performance might be sensitive to $\tau$ since it reflects the fixed number of data samples, but the results indicate that the performance is not significantly affected by variations in $\tau$. Moreover, EPQ algorithms with different $\tau$ consistently outperform the CQL baseline, highlighting the superiority of the proposed method.
@@ Appendix A Proof
where the equality in (A.4) satisfies when $\pi=\hat{\beta}$ or $f_{\tau}^{\pi,\hat{\beta}}=0$. Since all elements of $(I-\gamma P^{\pi})^{-1}$ are non-negative, $V_{\infty}$ underestimates the true value function $V^{\pi}$ if the penalizing constant $\alpha$ satisfies $\alpha\geq\max_{s,a\in D}[\xi^{\delta}(s,a)]\cdot\max_{s\in D}\Delta_{EPQ}^{\pi}(s)$. In addition, according to [15], the average penalty of CQL for policy actions can be represented as $\Delta_{CQL}^{\pi}(s)=\mathbb{E}_{a\sim\pi}[\frac{\pi}{\hat{\beta}}-1]$. Thus, $\Delta_{EPQ}^{\pi}(s)=f_{\tau}^{\pi,\hat{\beta}}(s)\Delta_{CQL}^{\pi}(s)$ and $f_{\tau}^{\pi,\hat{\beta}}(s)\leq 1$ from the definition in (2), so $0\leq\Delta_{EPQ}^{\pi}(s)\leq\Delta_{CQL}^{\pi}(s)$. In addition, if $\pi=\hat{\beta}$, then $0=\Delta_{EPQ}^{\hat{\beta}}(s)=\Delta_{CQL}^{\hat{\beta}}(s)$ from the equality condition in (A.4), which indicates that the average penalty for data actions is 00 for both EPQ and CQL. $\blacksquare$
@@ Appendix B Implementation Details
In this section, we provide the implementation details of the proposed EPQ. First of all, we provide a detailed derivation for the final $Q$-loss function (4) of EPQ in Section B.1. Next, we introduce a practical implementation of EPQ to compute the loss functions for the parameterized policy and $Q$-function in Section B.2. In addition, to calculate loss functions in Section B.2, we provide the additional implementation details in Appendices B.3, B.4, and B.5. We conduct our experiments on a single server equipped with an Intel Xeon Gold 6336Y CPU and one NVIDIA RTX A5000 GPU, and we compare the running time of EPQ with other baseline algorithms in Section B.6. For additional hyperparameters in the practical implementation of EPQ, we provide detailed hyperparameter setup and additional ablation studies in Appendix C and Appendix D, respectively.
@@ B.5 Implementation of $Q$-loss Function
$\log\sum_{a}\exp Q(s,a)=\log\left(\frac{1}{2}\sum_{a}\pi(a|s)\{\exp(Q(s,a)-\log\pi(a|s))\}+\frac{1}{2}\sum_{a}\rho\{\exp(Q(s,a)-\log\rho)\}\right)$
$\quad\quad\approx\log\left(\frac{1}{2N_{a}}\sum_{a_{n}\sim\pi}^{N_{a}}(\exp(Q(s,a_{n})-\log\pi(a_{n}|s)))+\frac{1}{2N_{a}}\sum_{a_{n}\sim\textrm{Unif}(\mathcal{A})}^{N_{a}}(\exp(Q(s,a_{n})-\log\rho)))\right),$ (B.7)
where $N_{a}$ is the number of action sampling and $\textrm{Unif}(\mathcal{A})$ is a Uniform distribution on $\mathcal{A}$ and $\rho$ is the density of uniform distribution.
@@ Appendix C Hyperparameter Setup
In addition, in Table 5, we provide the details of the task hyperparameters regarding our contributions in the proposed EPQ: the penalty control threshold $\tau$ and the IS clipping factor $c_{\min}$ in the $Q$-loss implementation in (B.2), and the cluster radius $\epsilon$ and regularizing temperature $\zeta$ for the practical implementation of IS clipping factor $w_{s,a}^{Q}$ in Section B.4. Note that $\rho$ in Table 5 represents the density of uniform distribution. For the task hyperparameters, we consider various hyperparameter setups and provide the best hyperparameter setup for all considered tasks in Table 5. The results are based on the ablations studies provided in Section 4.3 and Appendix D.
@@ Appendix D Additional Ablation Studies Related to $w_{s,a}^{Q}$ Estimation
IS Clipping Factor $c_{\min}$: In the EPQ implementation, the IS clipping factor $c_{\min}$ is employed to clip the IS weight $w_{s,a}^{Q}$ to prevent the exclusion of data samples with relatively low $w_{s,a}^{Q}$. When $c_{\min}=0$, low-quality samples with low $w_{s,a}^{Q}$ are not utilized at all based on the prioritization in Section 3.3. However, as $c_{\min}$ increases, these low-quality samples are increasingly exploited. Fig. LABEL:fig:ablation(c) illustrates the performance of EPQ with varying $c_{\min}$, and EPQ achieves the best performance when $c_{\min}=0.5$. This result suggests that it is more beneficial to use low-quality samples with proper priority rather than discarding them entirely.
@@ 33. Answer: [Yes]
  34. Justification: The graphs included in the paper such as Figure 5(a) and Figure LABEL:fig:ablation in Section 4 Experiments well demonstrate the statistical significance of the experiment.

### v2

@@ 4 Experiments
4.1 Performance Comparisons
We compare our algorithm with various constraint-based offline RL methods, including CQL baselines [21] on which our algorithm is based on. For other baseline methods, we consider behavior cloning (BC) and 10% BC, where the latter only utilizes only the top 10% of demonstrations with high returns, TD3+BC [27] that simply combines BC with TD3 [3], Onestep RL [28] that performs a single policy iteration based on the dataset, implicit $Q$-learning (IQL) [29] that seeks the optimal value function for the dataset through expectile regression, mildly conservative $Q$-learning (MCQ) [30] that reduces overestimation by using pseudo $Q$ values for out-of-distribution actions, and MISA [31] that considers the policy constraint based on mutual information. To assess baseline algorithm performance, we utilize results directly from the original papers for CQL (paper) [21] and MCQ [30], as well as reported results from other baseline algorithms according to Ma et al. [31]. For CQL, reproducing its performance is challenging, so we also include reproduced CQL performance labeled as CQL (reprod.) from Ma et al. [31]. Any missing experimental results have been filled in by re-implementing each baseline algorithm. For our algorithm, we explored various penalty control thresholds $\tau\in\{c\cdot\rho,~{}c\in[0,10]\}$, where $\rho$ represents the log-density of $\textrm{Unif}(\mathcal{A})$. For Mujoco tasks, the EPQ penalizing constant is fixed at $\alpha=20.0$, and for Adroit tasks, we consider either $\alpha=5.0$ or $\alpha=20.0$. To ensure robustness, we run our algorithm with four different seeds for each task. Table 1 displays the average normalized returns and corresponding standard deviations for compared algorithms. The performance of EPQ is based on the best hyperparameter setup, with additional results presented in the ablation study in Section 4.3. Further details on the hyperparameter setup are provided in Appendix C.
@@ 4.1 Performance Comparisons
4.2 The Analysis of Estimation Bias
In Section 4.1, EPQ outperforms CQL baselines significantly across various D4RL tasks based on our proposed penalty in Section 3. To analyze the impact of our overestimation reduction method on performance enhancement, we compare the estimation bias for EPQ and CQL baselines with various penalizing constants $\alpha\in\{0,1,5,10,20\}$ on ‘Hopper-random’, ‘Hopper-medium’, and ‘Halfcheetah-medium’ tasks. In Fig. 6(a), we depict the squared value of estimation bias, obtained from the difference between the $Q$-value and the empirical average return for sample trajectories generated by the policy, to show both overestimation bias and underestimation bias. In the experiment shown in Fig. 6(a), the estimation bias in CQL with $\alpha=0$ became excessively large, causing the gradients to explode and resulting in forced termination of the training. Fig. 6(b) illustrates the corresponding normalized average returns, emphasizing learning progress after $200\mathrm{k}$ gradient steps.
In Fig. 6(a), we observe an increase in estimation bias for CQL as the penalizing constant $\alpha$ rises, attributed to unnecessary bias highlighted in Fig. 1. Reducing $\alpha$ to nearly 0 in CQL, however, fails to effectively mitigate overestimation error, leading to a divergence of the $Q$-function in tasks such as ‘Hopper-random’ and ‘Hopper-medium’, as shown in Fig. 1. Conversely, EPQ demonstrates superior reduction of estimation bias in the $Q$-function compared to CQL baselines for all tasks in Fig. 6(a), indicating its capability to mitigate both overestimation and underestimation bias based on the proposed penalty. As a result, Fig. 6(b) shows that EPQ significantly outperforms all CQL variants on ‘Hopper-random’, ‘Hopper-medium’, and ‘Halfcheetah-medium’ tasks.
(a) Component evaluation
(b) Penalty control thresholds $\tau\in[0.2\rho,0.5\rho,1.0\rho,2.0\rho,5.0\rho,10.0\rho]$
Figure 7: Additional ablation studies on the Hopper-random, Hopper-medium, and Halfcheetah-medium tasks are presented. The best hyperparameter in the paper is denoted by the orange curve.
@@ 4.3 Ablation Study
To understand the impact of EPQ’s components and hyperparameters, we conduct ablation studies to evaluate each component and the penalty control threshold $\tau$ on the ‘Hopper-random’, ‘Hopper-medium’, and ‘HalfCheetah-medium’ tasks where our proposed method showed a significant performance improvement compared to the baseline CQL.
Component Evaluation: In Section 3, we introduced two variants of the EPQ algorithm: EPQ (w/o PD), which does not incorporate a prioritized dataset as in equation (3), and EPQ (with PD), which leverages a prioritized dataset based on $\hat{\beta}^{Q}$ as in equation (4). In Fig. 7(a), we compare the performance of EPQ (w/o PD), EPQ (with PD), and the CQL baseline to analyze the impact of each component. EPQ (w/o PD) still outperforms CQL, demonstrating that the proposed penalty $\mathcal{P}_{\tau}$ in Section 3.2 enhances performance by efficiently reducing overestimation without introducing unnecessary estimation bias, as discussed in Section 3.2. Additionally, Fig. 7(a) shows that EPQ (with PD) outperforms EPQ (w/o PD) significantly in the HalfCheetah-medium task, indicating that the proposed prioritized dataset contributes to improved performance, as anticipated in Section 3.3.
Penalty Control Threshold $\tau$: As discussed in Section 3.2, EPQ can dynamically control the penalty amount based on the penalty control threshold $\tau$, as illustrated in Fig. 2, even in the absence of knowledge about the exact number of data samples. Fig. 7(b) demonstrates the performance of EPQ with various penalty control thresholds $\tau\in[0.2\rho,0.5\rho,1.0\rho,2.0\rho,5.0\rho,10.0\rho]$, where $\rho$ represents the log-density of Unif$(\mathcal{A})$. Note that $\rho$ is negative, so $\tau=10.0\rho$ is the lowest threshold while $\tau=0.2\rho$ is the highest. The results indicate that in tasks like Hopper-medium, where a variety of actions are not sufficiently sampled, a higher threshold performs better. Conversely, in tasks like Hopper-random, where a broad range of actions is sampled, a lower threshold is more effective. An exception is the HalfCheetah-medium task, which, despite having fewer action variations, visits a diverse range of states. This may result in lower overestimation errors for OOD actions, benefiting from a lower threshold. Furthermore, the performance on the considered tasks appears to be surprisingly less sensitive to changes in $\tau$. We initially expect that performance might be sensitive to $\tau$ since it reflects the fixed number of data samples, but the results indicate that the performance is not significantly affected by variations in $\tau$. Moreover, EPQ algorithms with different $\tau$ consistently outperform the CQL baseline, highlighting the superiority of the proposed method.
@@ 6 Conclusion
Acknowledgements
This work was supported in part by Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.2022-0-00469, Development of Core Technologies for Task-oriented Reinforcement Learning for Commercialization of Autonomous Drones) and in part by IITP grant funded by the Korea government (MSIT) (No. RS-2022-00156361, Innovative Human Resource Development for Local Intellectualization(UNIST)) and in part by Artificial Intelligence Graduate School support (UNIST), IITP grant funded by the Korea government (MSIT) (No.2020-0-01336).
@@ Appendix A Proof
where the equality in (A.4) satisfies when $\pi=\hat{\beta}$ or $f_{\tau}^{\pi,\hat{\beta}}=0$. Given that $V_{k+1}$ converges to $V_{\infty}=V^{\pi}(s)+(I-\gamma P^{\pi})^{-1}\cdot\{-\alpha\Delta_{EPQ}^{\pi}(s)+\mathbb{E}_{a\sim\pi}[\xi^{\delta}(s,a)]\}$, choosing the penalizing constant $\alpha$ that satisfies $\alpha\geq\max_{s,a\in D}[\xi^{\delta}(s,a)]\cdot\max_{s\in D}(\Delta_{EPQ}^{\pi}(s))^{-1}$ will satisfy,
$-\alpha\cdot\Delta_{EPQ}^{\pi}(s)+\mathbb{E}_{a\sim\pi}[\xi^{\delta}(s,a)]$
$\leq-\max_{s,a\in D}[\xi^{\delta}(s,a)]\cdot\underbrace{\max_{s\in D}(\Delta_{EPQ}^{\pi}(s))^{-1}\cdot\Delta_{EPQ}^{\pi}(s)}_{\geq 1}+\mathbb{E}_{a\sim\pi}[\xi^{\delta}(s,a)]$
$\leq-\max_{s,a\in D}[\xi^{\delta}(s,a)]+\mathbb{E}_{a\sim\pi}[\xi^{\delta}(s,a)]\leq 0,~{}~{}~{}~{}\forall s,$ (A.5)
Since $I-\gamma P^{\pi}$ is non-singular $M$-matrix and the inverse of non-singular $M$-matrix is non-negative, i.e., all elements of $(I-\gamma P^{\pi})^{-1}$ are non-negative, $V_{\infty}(s)=V^{\pi}(s)+(I-\gamma P^{\pi})^{-1}\cdot\{-\alpha\Delta_{EPQ}^{\pi}(s)+\mathbb{E}_{a\sim\pi}[\xi^{\delta}(s,a)]\}\leq V^{\pi}(s),~{}\forall s.$ Therefore, $V_{\infty}$ underestimates the true value function $V^{\pi}$ if the penalizing constant $\alpha$ satisfies $\alpha\geq\max_{s,a\in D}[\xi^{\delta}(s,a)]\cdot\max_{s\in D}(\Delta_{EPQ}^{\pi}(s))^{-1}$. In addition, according to [21], the average penalty of CQL for policy actions can be represented as $\Delta_{CQL}^{\pi}(s)=\mathbb{E}_{a\sim\pi}[\frac{\pi}{\hat{\beta}}-1]$. Thus, $\Delta_{EPQ}^{\pi}(s)=f_{\tau}^{\pi,\hat{\beta}}(s)\Delta_{CQL}^{\pi}(s)$ and $f_{\tau}^{\pi,\hat{\beta}}(s)\leq 1$ from the definition in (2), so $0\leq\Delta_{EPQ}^{\pi}(s)\leq\Delta_{CQL}^{\pi}(s)$. In addition, if $\pi=\hat{\beta}$, then $0=\Delta_{EPQ}^{\hat{\beta}}(s)=\Delta_{CQL}^{\hat{\beta}}(s)$ from the equality condition in (A.4), which indicates that the average penalty for data actions is 00 for both EPQ and CQL. $\blacksquare$
@@ Appendix B Implementation Details
In this section, we provide the implementation details of the proposed EPQ. First of all, we provide a detailed derivation of the final $Q$-loss function(4) of EPQ in Section B.1. Next, we introduce a practical implementation of EPQ to compute the loss functions for the parameterized policy and $Q$-function in Section B.2. In addition, to calculate loss functions in Section B.2, we provide the additional implementation details in Appendices B.3, B.4, and B.5. We conduct our experiments on a single server equipped with an Intel Xeon Gold 6336Y CPU and one NVIDIA RTX A5000 GPU, and we compare the running time of EPQ with other baseline algorithms in Section B.6. For additional hyperparameters in the practical implementation of EPQ, we provide detailed hyperparameter setup and additional ablation studies in Appendix C and Appendix D, respectively.
@@ B.5 Implementation of $Q$-loss Function
$\log\sum_{a}\exp Q(s,a)=\log\left(\frac{1}{2}\sum_{a}\pi(a|s)\{\exp(Q(s,a)-\log\pi(a|s))\}+\frac{1}{2}\sum_{a}\rho_{d}\{\exp(Q(s,a)-\log\rho_{d})\}\right)$
$\quad\quad\approx\log\left(\frac{1}{2N_{a}}\sum_{a_{n}\sim\pi}^{N_{a}}(\exp(Q(s,a_{n})-\log\pi(a_{n}|s)))+\frac{1}{2N_{a}}\sum_{a_{n}\sim\textrm{Unif}(\mathcal{A})}^{N_{a}}(\exp(Q(s,a_{n})-\log\rho_{d})))\right),$ (B.7)
where $N_{a}$ is the number of action sampling, $\textrm{Unif}(\mathcal{A})$ is a Uniform distribution on $\mathcal{A}$, and $\rho_{d}$ is the density of uniform distribution.
@@ Appendix C Hyperparameter Setup
In addition, in Table 5, we provide the details of the task hyperparameters regarding our contributions in the proposed EPQ: the penalty control threshold $\tau$ and the IS clipping factor $c_{\min}$ in the $Q$-loss implementation in (B.2), and the cluster radius $\epsilon$ and regularizing temperature $\zeta$ for the practical implementation of IS clipping factor $w_{s,a}^{Q}$ in Section B.4. Note that $\rho$ in Table 5 represents the log-density of uniform distribution. For the task hyperparameters, we consider various hyperparameter setups and provide the best hyperparameter setup for all considered tasks in Table 5. The results are based on the ablations studies provided in Section 4.3 and Appendix D.
@@ Appendix D Additional Ablation Studies Related to $w_{s,a}^{Q}$ Estimation
IS Clipping Factor $c_{\min}$: In the EPQ implementation, the IS clipping factor $c_{\min}$ is employed to clip the IS weight $w_{s,a}^{Q}$ to prevent the exclusion of data samples with relatively low $w_{s,a}^{Q}$. When $c_{\min}=0$, low-quality samples with low $w_{s,a}^{Q}$ are not utilized at all based on the prioritization in Section 3.3. However, as $c_{\min}$ increases, these low-quality samples are increasingly exploited. Fig. 7(c) illustrates the performance of EPQ with varying $c_{\min}$, and EPQ achieves the best performance when $c_{\min}=0.5$. This result suggests that it is more beneficial to use low-quality samples with proper priority rather than discarding them entirely.
@@ 33. Answer: [Yes]
  34. Justification: The graphs included in the paper such as Figure 6 and Figure 7 in Section 4 Experiments well demonstrate the statistical significance of the experiment.